OpenStudy (anonymous):

what does e=mc3 mean?

HanAkoSolo (jamierox4ev3r):

it is an equation by Albert Einstein showing that energy and mass are interchangeable. E is energy. M is mass. C is the speed of light thus energy. DOes that make sense @daniellecrook ?

OpenStudy (anonymous):

i think you mean \[\epsilon=mc^2\]

HanAkoSolo (jamierox4ev3r):

and yes it is \(E=mc^2\)

OpenStudy (anonymous):

not really can you explain a little bit more? @Jamierox4ev3r

OpenStudy (anonymous):

It is the equation that helped prove that matter and energy are not separate but different forms of the same thing. It means the energy contained in matter (e) is equal to the mass of the matter (m) times the speed of light (c) squared.

OpenStudy (anonymous):

i dont understand math im so bad at it :(

HanAkoSolo (jamierox4ev3r):

it is an equation that energy and mass are basically undistinguishable. This formula is also often used to depict energy.

HanAkoSolo (jamierox4ev3r):

@ivanapen8 is you have any questions, I'd be glad to help you :)

OpenStudy (anonymous):

oh ok i get it now thank you i will vote you best response!

OpenStudy (anonymous):

@Jamierox4ev3r aw thank you

HanAkoSolo (jamierox4ev3r):

no problem, welcome to Openstudy!

OpenStudy (anonymous):

@Jamierox4ev3r How is the speed of light = energy?

HanAkoSolo (jamierox4ev3r):

well light, regardless of the forms, produces energy and often times heat as a by-product.

OpenStudy (anonymous):

If what you are saying is true based on the foundations and laws of physics then \(\bf E = c\) should suffice. We don't even the mass and there is no need to square the speed of light either. @jamierox4ev3r

OpenStudy (anonymous):

@jamierox4ev3r your explanation leaves much to be desired. @daniellecrook it is an equation from Einstein's theory of special relativity that relates the mass of a body at rest to its rest energy. It is a specialized form of the broader energy-momentum relation:$$E^2=(pc)^2+(m_0c^2)^2$$

OpenStudy (anonymous):

@oldrin.bataku Did you just look at that up?

OpenStudy (anonymous):

@genius12 no hes smart af so u can leave

OpenStudy (anonymous):

sorry @jamierox3ev3r i gotta give best reponse to @oldrin.bataku

OpenStudy (anonymous):

@ivanapen8 Sorry sir but I'm here to learn. It's best if you do the same too.

OpenStudy (anonymous):

@genius12 no it's taught in a basic high school physics class that touches on special relativity. It is more elementary than Lorentz and Poincare groups, that's for sure.

OpenStudy (anonymous):

It's taught at around the same time the Lorentz factor is introduced to explain time dilation, length contraction, etc. for bodies moving at relativistic speeds in inertial reference frames

OpenStudy (anonymous):

What's p in the equation?

OpenStudy (anonymous):

As stated in its title, it is the momentum of the body.

OpenStudy (anonymous):

I know what quantity does p represent?

OpenStudy (anonymous):

tbh this was a joke post @genius12 you can leave

OpenStudy (anonymous):

I just told you it is the (scalar) momentum of the body.

OpenStudy (anonymous):

@oldrin.bataku How is momentum a scalar quantity?

OpenStudy (anonymous):

clearly it is the magnitude of the momentum vector